OpenStudy (anonymous):

Need help explaining how this would be completed. --------------------------------- Landy made a gift box in the size shown below. He wants to stick a ribbon along the diagonal of the top face of the box. The length of the piece of ribbon he has is 20 inches. Is the length of the piece of ribbon long enough to be placed on the diagonal of the top?

OpenStudy (anonymous):

A. Yes, because \[\left(\begin{matrix}18+15 \\ 2\end{matrix}\right)\] is smaller than 20.

OpenStudy (anonymous):

Heres the pic btw

OpenStudy (anonymous):

B. No, because \[\sqrt{2\left( 18^{2}+15^{2} \right)}\] is greater than 20. ---------------------- C. Yes, because \[\sqrt{18^{2}}-15^{2}\]is smaller than 20. ---------------------- D. No, because \[\sqrt{18^{2}}+15^{2}\] is greater than 20.

OpenStudy (accessdenied):

 It's a right triangle, so Pythagorean Theorem must apply here. Logically, for a \(20\) inch ribbon to be long enough to reach across that diagonal, the length of the diagonal should be less than or equal to \(20\).